The average angular acceleration of a s tone lodged in a rotating car wheel is 100 rads -2 what will the stones final angular speed be after 3.0 s if it starts from rest ? Also calculate it's angular displacement during this time
1
Expert's answer
2021-10-28T08:53:44-0400

Explanations& Calculations


  • Assuming the stone would remain in contact with the tyre without dislodging from it,

a)

  • Apply "\small \omega_f = \omega_i +\alpha t" for the motion, you can compute the final angular velocity.

"\qquad\qquad
\begin{aligned}
\small\omega_f&=\small 0+100\,rads^{-2}\times3.0\,s\\
&=\small \bold{300\,rads^{-1}}
\end{aligned}"

b)

  • Apply "\small \theta= \omega_it+\frac{1}{2}\alpha t^2" for the motion, you can compute the angular displacement of the stone.

"\qquad\qquad
\begin{aligned}
\small \theta&=\small 0\times3.0\,s+0.5\times(100\,rads^{-2})\times(3.0\,s)^2\\
&=\small \bold{450\,rad}
\end{aligned}"

  • This means the stone has completed a full rotation ("\small 2\pi \,rad = 360^0" ) and another quarter ("\small \pi/2\,rad =90^0") during this 3 seconds.